OpenStudy (anonymous):

Fx=F(x)-F(a)/(x-a) where Fx= x^2-5 and the point is (3,4)

OpenStudy (anonymous):

So i know my equation ends up being: (X^2-5)-(4)/(x-3) = (x^2-9)/ (x-3) = (x-3)(x+3)/(x-3) = x+3/1 My only question is to why the x is another 3 and my answer is 6? shouldn't x=4?

OpenStudy (amoodarya):

in point (3,4) x=3 , y=4 so f(3)=9-5=4 you compute the formula by 'x" not by "y"